OpenStudy (anonymous):

solve each system of equations using elimination or substitution: 2x-3y=8 x+y=4 please, please, PLEASE help me... I am so confuzzled its scary!

OpenStudy (anonymous):

Lets do substitution. Step 1: Solve x+y=4 for x. Can you do that?

OpenStudy (anonymous):

soory was having lunch. no i do not know how to do that, but i am a fast learner and if you coud just explain i would be so happy

OpenStudy (texaschic101):

x + y = 4.....we want x on one side everything else on the other side, so subtract y from both sides and what do you get ?

OpenStudy (texaschic101):

oh...the asker just left......he/she was just here

OpenStudy (texaschic101):

must not want to learn, just wants the answer

OpenStudy (anonymous):

no no no no no.... i was just checking on my baby sister. she fell and needed a bandaid. please stay i really do want to learn and no i dont just want the answer

OpenStudy (anonymous):

x=4-y right?

OpenStudy (anonymous):

please stay with me. i just need to figure this out :(

OpenStudy (texaschic101):

yes....x = 4 - y now that we know what x is we substitute 4 - y in for x in the other equation. can you do that ?

OpenStudy (anonymous):

so like 2(4-y)-3y=8?

OpenStudy (anonymous):

8-2y-3y=8

OpenStudy (texaschic101):

yes...now combine like terms

OpenStudy (texaschic101):

you need help

OpenStudy (anonymous):

how do i combine terms? hold on i need to get my charger.. this will take like to seconds pease dont disappear my laptop is about to die on me:)

OpenStudy (texaschic101):

add your y's together

OpenStudy (anonymous):

ok im back... so how do i combine terms

OpenStudy (texaschic101):

add the y's together

OpenStudy (anonymous):

2y +3y=5y?

OpenStudy (anonymous):

or is it -2y+3y=1y?

OpenStudy (texaschic101):

no... 8 - 2y - 3y = 8 (add the y's) 8 - 5y = 8 now we want 5y by itself so subtract 8 from both sides -5y = 8-8 -5y = 0 y = 0/-5 which is 0 do you understand ?

OpenStudy (anonymous):

yeah, i am following :)

OpenStudy (texaschic101):

so we know that y = 0. Now sub 0 in for y in either of your original equations to find x

OpenStudy (anonymous):

so 2(4-0)-0=8? 8-0=8???

OpenStudy (texaschic101):

2x - 3y = 8 2x - 3(0) = 8 2x - 0 = 8 2x = 8 x = 8/2 x = 4 do you understand ? or you could have subbed it into the other equation: x + y = 4 x + 0 = 4 x = 4

OpenStudy (anonymous):

wait!!! okay i think i figured it out. y=0 and x=4 once i put it as an ordered pair it makes complet and total sense!!!

OpenStudy (texaschic101):

you can check your answers by subbing in known values into either or both or your original equations... 2x + 3y = 8 2(4) + 3(0) = 8 8 + 0 = 8 8 = 8 (correct) if the final result comes out equal, then you have the correct answers

OpenStudy (anonymous):

ok thank you... i gotta go clean my room so bye-bye. thanks for the help

OpenStudy (texaschic101):

your welcome.....sorry about what I said earlier....I was having a bad morning

OpenStudy (texaschic101):

Is your baby sister ok ?

OpenStudy (texaschic101):

I am sure she is....have fun cleaning your room....make it a game, it makes it easier

OpenStudy (anonymous):

thanks. you are like my math super hero right now and yeah its totally cool. i rage on people randomly all the time and yeah my sisters fine thanks bye